# Exploring Inevitable Waypoints for Unsolvability Explanation in Hybrid Planning Problems

### v1

@@ Abstract.
Explainable AI Planning and Hybrid Systems and Unsolvability and Sub-problem and Bounded Reachability.
@@
ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interactive systems and toolsccs: Theory of computation Timed and hybrid models
@@ 1. Introduction
In human-computer interaction (HCI), humans engage in trustworthy collaborations with autonomous agents, and one of the precursors of such collaborations is that an autonomous agent must explain the rationale behind its decision to the human. With the emergence of artificial intelligence (AI) and the multitude of application domains where AI planning is being envisioned to replace plans generated by humans, Explainable AI Planning (XAIP) (DBLP:journals/corr/abs-1709-10256, ; DBLP:conf/rweb/HoffmannM19, ) has emerged as an important connection between HCI and AI for designing explainable systems that bridges the gap between theoretical and algorithmic planning and real-world applications (ijcai2020p669, ). While there has been a lot of research on generating explanations to planning problems, most of the earlier works in explanation generation have focused on explaining why a given plan or action was chosen (DBLP:conf/ijcai/ChakrabortiSZK17, ; DBLP:conf/aips/ChakrabortiKSSK19, ; DBLP:journals/jair/KrarupKMLC021, ; DBLP:journals/tecs/SarwarRB23, ). However, explaining the unsolvability of a given planning problem remains a largely open and understudied problem. The recent works that focus on explaining the unsolvability of planning problems have primarily concentrated on generating certificates or proofs of unsolvability (DBLP:conf/aips/ErikssonRH17, ; DBLP:conf/aips/ErikssonRH18, ), or on identifying counterfactual alterations to the original planning task to make it solvable, often referred to as “excuses” (DBLP:conf/aips/GobelbeckerKEBN10, ). These approaches, which are more oriented towards automatic verification, may fall short in adequately explaining unsolvability in complex planning domains.
@@
A well-known insight into human thinking and problem-solving is that humans tend to decompose a problem into sub-problems that help in progressively converging towards the goal. Many AI systems mimic this notion in the way they solve problems. For instance, the main feature of the pioneering automated theorem prover, logic theorist, is the use of problem-subproblem hierarchy (LT-1956, ). An innovative technique for identification of sub-problems relevant for explaining unsolvability of a planning problem in domains with discrete dynamics has been proposed in (DBLP:conf/ijcai/SreedharanSSK19, ). In this paper, we propose to adopt the same philosophy of sub-problem identification as an efficient mechanism for analyzing and explaining unsolvability of planning problems in hybrid domains, domains with a combination of discrete and continuous dynamics. In particular, for a given unsolvable planning problem, we propose to identify a sequence of waypoints, which are universal obstacles to plan existence, in other words, they appear on every path on every plan from the source to the planning goal. This work envisions such waypoints as sub-problems of the planning problem and the unreachability of any of these waypoints as an explanation for the unsolvability of the original problem at hand. We propose a novel method of waypoint identification by casting the problem as an instance of the longest common subsequence problem, a widely popular problem in computer science, typically considered as an illustrative example for the dynamic programming paradigm. Once the waypoints are identified, we perform symbolic reachability analysis on them to identify the earliest unreachable waypoint and report it as the explanation of unsolvability. We present experimental results on unsolvable planning problems in hybrid domains. In summary, the key contributions of this paper are:
@@ 2. Motivating Example
In this section, we present a motivating example in the context of a planning problem for a planetary rover that explores a planetary site and collects samples for experiments. The agent (autonomous battery powered rover) possesses knowledge about the topography of the exploration site as a planar grid. Figure 1 shows the topography as a 5$\times$5 grid of cells. The rover is initially positioned at cell 11 and there is a base-station at cell 25. The task of the rover is to collect soil and rock samples from designated sites, in cells 1 and 14 respectively and then reach the base-station. Mountainous regions and craters in the terrain are marked as impassable (cells 7, 12, etc). There are inclined areas in the terrain shown in Orange cells. The motion dynamics of the rover is interpreted as a hybrid system. The rover’s continuous motion and battery discharge dynamics can be different in each cell. For instance, the motion and battery depletion dynamics in an inclined region is different from the dynamics in flat-regions and regions where soil and rock samples are collected. When a rover makes a transition from one cell to another, it starts to follow the dynamics of the new cell instantaneously. The discrete dynamics here capture the connectivity of the cells in the presence of mountains and craters where the rover cannot move. The rover’s movement is restricted to one of its adjacent cells and movement to diagonal cells is prohibited.
@@
When a planner reports the planning task as unsolvable, our algorithm identifies ordered waypoints that ought to be reached in any plan in order to achieve the task. For instance, in the discussed domain, our algorithm detects that the cells 6-1-2-3-8-13-14 must be visited by any valid plan to solve the planning task. These are marked as ordered waypoints $w1-w7$ in the figure. Our proposed algorithm envisions such waypoints as sub-goals of the planning problem. The unreachability of any of these waypoints under the domain dynamics is reported as an explanation to the unsolvability of the original planning problem. The computational challenge lies in finding the waypoints, finding an order between them and lastly, finding the earliest waypoint which is unreachable under the dynamics. We propose a method of waypoint identification by casting the problem as an instance of the longest common subsequence problem, a widely popular problem in computer science, typically considered as an illustrative example for the dynamic programming paradigm. Unreachability of a waypoint is determined using a bounded model checker. For instance, given the initial rover battery charge of 10 units and the battery depletion rates of the cells (depletion rate of 1 unit in cells except in cell 1 and cell 14 where soil and rock sampling depletes battery at a higher rate of 2 units, and in the cells with inclination having a depletion rate of 3 units), reachability analysis determines that cell 13 is the first unreachable waypoint and reports this as an explanation of unsolvability.
Some of the waypoints as sub-goals may be explicitly known from the planning problem description itself. For example, the planetary rover domain has two sub-goals explicitly mentioned, namely the collection of soil and rock samples from cell 1 and cell 14 respectively, marked as waypoints $w2$ and $w7$ respectively. There may be sub-goals that are not apparent from the problem description explicitly but they are implicitly mandatory to complete the bigger planning task. We term these as inevitable waypoints. For example, the inevitable waypoints in our planetary rover domain are marked $w1$ and $w3-w6$ in the figure. Our waypoint detection algorithm detects both the explicit as well as implicit waypoints. In the following section, we formally describe the domain representation and the explanation problem we intend to solve.
@@ 3. Problem Overview
We begin with a formal definition of a planning problem in a hybrid system. A hybrid system exhibits an interplay of discrete and continuous dynamics. Hybrid automata are a well-known mathematical model for such systems (ALUR19953, ; 10.1007/3-540-57318-6_30, ). We now define a hybrid automaton model of a hybrid system.
@@
(ALUR19953, ) A hybrid automaton (HA) is a seven-tuple HA=(Loc, Var, Flow, Init, Lab, Edge, Inv) where:
@@
[…]
@@
[…]
@@
[…]
  • $Dom$ (DBLP:journals/tecs/SarwarRB23, ) is represented as a hybrid automaton HA.
  • $Prob$ is a tuple $\langle Init,Goal\rangle$ representing a problem description, where $Init$ and $Goal$ define the initial and the goal states.
@@
A plan for a planning problem $\Pi$ is a tuple $\langle$ $\lambda_{n}$, $makespan$rangle where $\lambda_{n}$ is a finite sequence of $n$ pairs $\langle t_{i},a_{i}\rangle$. In the pair, $t_{i}\in\mathbb{R}^{+}$ is the time instance of executing the action $a_{i}\in$ Lab. In the sequence, $t_{i}$ is non-decreasing. The $makespan$ is the duration of the plan. $\Box$
@@ 4.1. Decomposition into Sub-Problems
Given planning problems $\Pi_{i}$ = $(Dom_{i}$, $Prob_{i}$, $Depth_{i})$ and $\Pi_{j}$ = $(Dom_{j}$, $Prob_{j}$, $Depth_{j})$, we say $\Pi_{i}$ is a sub-problem of $\Pi_{j}$ when the following conditions hold: (a) $Dom_{i}=Dom_{j}$, (b) $Depth_{i}=Depth_{j}$, and (c) $Prob_{i}$, $Prob_{j}$ are tuples $\langle$ $Init_{i}$, $Goal_{i}$rangle and $\langle$ $Init_{j}$, $Goal_{j}$rangle resp. where
@@ 4.1.1. Abstraction over Planning Problems
where Loc is the finite set of locations of HA, $\Pi_{\ell}$ = $\langle$ $HA$, $\langle Init$, $Goal_{\ell}\rangle$, $Depth$rangle and $Goal_{\ell}=\langle\ell,Inv(\ell)\rangle$.
@@
For $\langle\Pi^{*},<\rangle$ to be poset, the binary relation $<$ on $\Pi^{*}$ should be reflexive, anti-symmetric, and transitive, that is, $<$ must be a partial order relation. From definition 4.1, it is easy to see that every planning problem in $\Pi^{*}$ is a sub-problem to itself and hence reflexive. Transitivity and anti-symmetry also follow from the definition of sub-problem. ∎
@@
The explanation is thus the first unsolvable sub-problem of $\Pi$ in a chain of sub-problems in the poset $\langle\Pi^{*},<\rangle$. As an illustration, assume that $Y^{<}=\{\Pi_{1}$, $\Pi_{2}$, $\ldots$, $\Pi_{n}\}$ is a chain of sub-problems of $\Pi$ in $\langle\Pi^{*},<\rangle$ having a total order as $\Pi_{1}<\Pi_{2}<\ldots<\Pi_{n}$, Explanation($\Pi$) is the $\Pi_{i}\in Y^{<}$ that is unsolvable where $\forall\Pi_{j}\in Y^{<}$ such that $\Pi_{j}<\Pi_{i}$, $\Pi_{j}$ is solvable. For instance, consider the planning problem of the motivating example. Our abstraction will render 25 planning problems in $\Pi^{*}$, each having one of the cell as the goal. An example of a chain of sub-problems is the chain consisting of 7 sub-problems shown as waypoints $w1-w7$ with a total order $w1<w2<\ldots<w7$ amongst them. Observe that this chain is also the longest possible chain of sub-problems in the poset. In this chain, the explanation generated will be the earliest $w_{i}$ that is unsolvable.
@@ 4.2. Reduction to Longest Common Subsequence (LCS) Problem
[…]
@@
A path $p$ between locations $l_{0}$ and $l_{n}$ in $\mathcal{G_{HA}}$ is a sequence of $locations$ and $edges$ given as:
@@
where $l_{i}$in $Loc$, $e_{i}$in $Edge$, and $l_{i}$, $l_{i+1}$ are the source and destination of $e_{i}$ respectively. The length of a path is the number of edges it contains. $\Box$
@@
The graph of the hybrid automaton provides a higher abstraction of the domain in the sense that if there is no path from $l_{0}$ to $l_{goal}$ in $\mathcal{G_{HA}}$, then there cannot be any valid run of a plan from $Init$ to $Goal$ and hence the planning problem is unsolvable. We may then identify the cause of unsolvability to be in the discrete dynamics, oblivious to the continuous dynamics of the domain. More importantly, as we shall see now, a chain of sub-problems can be identified from the longest common subsequence of the strings in $PS(\Pi)$. Finding a longest common subsequence (LCS) between strings is a classic computer science problem. An LCS measures the closeness of two or more strings by finding the maximum number of identical symbols in them in the same order (DBLP:conf/spire/BergrothHR00, ; DBLP:journals/jacm/Maier78, ; princeton1974finding, ). Recall that a subsequence is different from a substring which additionally requires that the common symbols present in the strings are without gaps. We now present the main result of the paper.
@@
LCS of strings can be computed using the standard dynamic programming paradigm (DBLP:conf/spire/BergrothHR00, ). For instance, in the motivating example of planetary rover domain, the longest common sequence for the strings of paths of length bounded by 10 from $l_{11}$ to $l_{25}$ is $l_{11}$-$l_{6}$-$l_{1}$-$l_{2}$-$l_{3}$-$l_{8}$-$l_{13}$-$l_{14}$-$l_{25}$. In the next section, we show the computation of Explanation($\Pi$).

### v2

@@ Abstract.
$c$s: Computing methodologies Artificial intelligenceccs: Human-centered computing Interactive systems and toolsccs: Theory of computation Timed and hybrid models
@@ 1. Introduction
In human-computer interaction (HCI), humans engage in trustworthy collaborations with autonomous agents, and one of the precursors of such collaborations is that an autonomous agent must explain the rationale behind its decision to the human. With the emergence of artificial intelligence (AI) and the multitude of application domains where AI planning is being envisioned to replace plans generated by humans, Explainable AI Planning (XAIP) (DBLP:journals/corr/abs-1709-10256; DBLP:conf/rweb/HoffmannM19) has emerged as an important connection between HCI and AI for designing explainable systems that bridges the gap between theoretical and algorithmic planning and real-world applications (ijcai2020p669). While there has been a lot of research on generating explanations to planning problems, most of the earlier works in explanation generation have focused on explaining why a given plan or action was chosen (DBLP:conf/ijcai/ChakrabortiSZK17; DBLP:conf/aips/ChakrabortiKSSK19; DBLP:journals/jair/KrarupKMLC021; DBLP:journals/tecs/SarwarRB23). However, explaining the unsolvability of a given planning problem remains a largely open and understudied problem. The recent works that focus on explaining the unsolvability of planning problems have primarily concentrated on generating certificates or proofs of unsolvability (DBLP:conf/aips/ErikssonRH17; DBLP:conf/aips/ErikssonRH18), or on identifying counterfactual alterations to the original planning task to make it solvable, often referred to as “excuses” (DBLP:conf/aips/GobelbeckerKEBN10). These approaches, which are more oriented towards automatic verification, may fall short in adequately explaining unsolvability in complex planning domains.
@@
A well-known insight into human thinking and problem-solving is that humans tend to decompose a problem into sub-problems that help in progressively converging towards the goal. Many AI systems mimic this notion in the way they solve problems. For instance, the main feature of the pioneering automated theorem prover, logic theorist, is the use of problem-subproblem hierarchy (LT-1956). An innovative technique for identification of sub-problems relevant for explaining unsolvability of a planning problem in domains with discrete dynamics has been proposed in (DBLP:conf/ijcai/SreedharanSSK19). In this paper, we propose to adopt the same philosophy of sub-problem identification as an efficient mechanism for analyzing and explaining unsolvability of planning problems in hybrid domains, domains with a combination of discrete and continuous dynamics. In particular, for a given unsolvable planning problem, we propose to identify a sequence of waypoints, which are universal obstacles to plan existence, in other words, they appear on every path on every plan from the source to the planning goal. This work envisions such waypoints as sub-problems of the planning problem and the unreachability of any of these waypoints as an explanation for the unsolvability of the original problem at hand. We propose a novel method of waypoint identification by casting the problem as an instance of the longest common subsequence problem, a widely popular problem in computer science, typically considered as an illustrative example for the dynamic programming paradigm. Once the waypoints are identified, we perform symbolic reachability analysis on them to identify the earliest unreachable waypoint and report it as the explanation of unsolvability. We present experimental results on unsolvable planning problems in hybrid domains. In summary, the key contributions of this paper are:
@@ 2. Motivating Example
In this section, we present a motivating example in the context of a planning problem for a planetary rover that explores a planetary site and collects samples for experiments. The agent (autonomous battery powered rover) possesses knowledge about the topography of the exploration site as a planar grid. Figure 1 shows the topography as a $\times5$ grid of cells. The rover is initially positioned at cell 11 and there is a base-station at cell 25. The task of the rover is to collect soil and rock samples from designated sites, in cells 1 and 14 respectively and then reach the base-station. Mountainous regions and craters in the terrain are marked as impassable (cells 7, 12, etc). There are inclined areas in the terrain shown in Orange cells. The motion dynamics of the rover is interpreted as a hybrid system. The rover’s continuous motion and battery discharge dynamics can be different in each cell. For instance, the motion and battery depletion dynamics in an inclined region is different from the dynamics in flat-regions and regions where soil and rock samples are collected. When a rover makes a transition from one cell to another, it starts to follow the dynamics of the new cell instantaneously. The discrete dynamics here capture the connectivity of the cells in the presence of mountains and craters where the rover cannot move. The rover’s movement is restricted to one of its adjacent cells and movement to diagonal cells is prohibited.
@@
When a planner reports the planning task as unsolvable, our algorithm identifies ordered waypoints that ought to be reached in any plan in order to achieve the task. For instance, in the discussed domain, our algorithm detects that the cells 6-1-2-3-8-13-14 must be visited by any valid plan to solve the planning task. These are marked as ordered waypoints w1w7w1-w7 in the figure. Our proposed algorithm envisions such waypoints as sub-goals of the planning problem. The unreachability of any of these waypoints under the domain dynamics is reported as an explanation to the unsolvability of the original planning problem. The computational challenge lies in finding the waypoints, finding an order between them and lastly, finding the earliest waypoint which is unreachable under the dynamics. We propose a method of waypoint identification by casting the problem as an instance of the longest common subsequence problem, a widely popular problem in computer science, typically considered as an illustrative example for the dynamic programming paradigm. Unreachability of a waypoint is determined using a bounded model checker. For instance, given the initial rover battery charge of 10 units and the battery depletion rates of the cells (depletion rate of 1 unit in cells except in cell 1 and cell 14 where soil and rock sampling depletes battery at a higher rate of 2 units, and in the cells with inclination having a depletion rate of 3 units), reachability analysis determines that cell 13 is the first unreachable waypoint and reports this as an explanation of unsolvability.
Some of the waypoints as sub-goals may be explicitly known from the planning problem description itself. For example, the planetary rover domain has two sub-goals explicitly mentioned, namely the collection of soil and rock samples from cell 1 and cell 14 respectively, marked as waypoints $w2$ and $w7$ respectively. There may be sub-goals that are not apparent from the problem description explicitly but they are implicitly mandatory to complete the bigger planning task. We term these as inevitable waypoints. For example, the inevitable waypoints in our planetary rover domain are marked $w1$ and w3w6w3-w6 in the figure. Our waypoint detection algorithm detects both the explicit as well as implicit waypoints. In the following section, we formally describe the domain representation and the explanation problem we intend to solve.
@@ 3. Problem Overview
We begin with a formal definition of a planning problem in a hybrid system. A hybrid system exhibits an interplay of discrete and continuous dynamics. Hybrid automata are a well-known mathematical model for such systems (ALUR19953; 10.1007/3-540-57318-6_30). We now define a hybrid automaton model of a hybrid system.
@@
(ALUR19953) A hybrid automaton (HA) is a seven-tuple HA=(Loc, Var, Flow, Init, Lab, Edge, Inv) where:
@@
  • Flow is a mapping from each location $l\in\emph{Loc}$ to a set of differential equations $\{\dot{x}\in f(x_{1},\ldots,x_{|Var|})\mid x\in\emph{Var}\}$, where $\dot{x}$ denotes the rate of change of variable $x$.
  • Init is a tuple $\langle l_{0},S\rangle$ such that $l_{0}\in\emph{Loc}$ and $S\subseteq V$.
@@
  • Edge is a finite set of transitions e=(l,a,g,r,l)e=(l,a,g,r,l’), each consisting of a source location $l\in\emph{Loc}$, a target location $l^{\prime}\in\emph{Loc}$, a label $a\in\emph{Lab}$, a guard $g\subseteq V$ and a reset map $r:\mathbb{R}^{|Var|}\to 2^{\mathbb{R}^{|Var|}}$.
  • Inv is a mapping from each location $l\in\emph{Loc}$ to a subset of valuations $V$. $\Box$
@@
A planning problem $\Pi$ for a hybrid system is a three-tuple $(Dom$, ProbProb, $Depth)$, where
  • DomDom (DBLP:journals/tecs/SarwarRB23) is represented as a hybrid automaton HA.
  • ProbProb is a tuple $\langle Init,Goal\rangle$ representing a problem description, where $Init$ and $Goal$ define the initial and the goal states.
@@
A plan for a planning problem $\Pi$ is a tuple $\langle$ $\lambda_{n}$, makespanmakespan $\rangle$ where $\lambda_{n}$ is a finite sequence of $n$ pairs $\langle t_{i},a_{i}\rangle$. In the pair, $t_{i}\in\mathbb{R}^{+}$ is the time instance of executing the action $a_{i}\in$ Lab. In the sequence, $t_{i}$ is non-decreasing. The makespanmakespan is the duration of the plan. $\Box$
@@ 4.1. Decomposition into Sub-Problems
Given planning problems $\Pi_{i}$ = $(Dom_{i}$, $Prob_{i}$, $Depth_{i})$ and $\Pi_{j}$ = $(Dom_{j}$, $Prob_{j}$, $Depth_{j})$, we say $\Pi_{i}$ is a sub-problem of $\Pi_{j}$ when the following conditions hold: (a) $Dom_{i}=Dom_{j}$, (b) $Depth_{i}=Depth_{j}$, and (c) $Prob_{i}$, $Prob_{j}$ are tuples $\langle$ $Init_{i}$, $Goal_{i}$ $\rangle$ and $\langle$ $Init_{j}$, $Goal_{j}$ $\rangle$ resp. where
@@ 4.1.1. Abstraction over Planning Problems
where Loc is the finite set of locations of HA, $\Pi_{\ell}$ = $\langle$ HAHA, $\langle Init$, $Goal_{\ell}\rangle$, DepthDepth $\rangle$ and $Goal_{\ell}=\langle\ell,Inv(\ell)\rangle$.
@@
For $\langle\Pi^{*},<\rangle$ to be poset, the binary relation << on $\Pi^{*}$ should be reflexive, anti-symmetric, and transitive, that is, << must be a partial order relation. From definition 4.1, it is easy to see that every planning problem in $\Pi^{*}$ is a sub-problem to itself and hence reflexive. Transitivity and anti-symmetry also follow from the definition of sub-problem. ∎
@@
The explanation is thus the first unsolvable sub-problem of $\Pi$ in a chain of sub-problems in the poset $\langle\Pi^{*},<\rangle$. As an illustration, assume that $Y^{<}=\{\Pi_{1}$, $\Pi_{2}$, $\ldots$, $\Pi_{n}\}$ is a chain of sub-problems of $\Pi$ in $\langle\Pi^{*},<\rangle$ having a total order as $\Pi_{1}<\Pi_{2}<\ldots<\Pi_{n}$, Explanation($\Pi$) is the $\Pi_{i}\in Y^{<}$ that is unsolvable where $\forall\Pi_{j}\in Y^{<}$ such that $\Pi_{j}<\Pi_{i}$, $\Pi_{j}$ is solvable. For instance, consider the planning problem of the motivating example. Our abstraction will render 25 planning problems in $\Pi^{*}$, each having one of the cell as the goal. An example of a chain of sub-problems is the chain consisting of 7 sub-problems shown as waypoints w1w7w1-w7 with a total order $w1<w2<\ldots<w7$ amongst them. Observe that this chain is also the longest possible chain of sub-problems in the poset. In this chain, the explanation generated will be the earliest $w_{i}$ that is unsolvable.
@@ 4.2. Reduction to Longest Common Subsequence (LCS) Problem
The graph of a hybrid automaton HA is defined as $\mathcal{G_{HA}}=(V$, $E)$, where $V$ = Loc and $E\subseteq\emph{Loc}\times\emph{Loc}$ such that for every $(l,a,g,r,l^{\prime})\in\emph{Edge}$, there is an edge $(l,l^{\prime})\in E$. $\Box$
@@
A path $p$ between locations $l_{0}$ and $l_{n}$ in $\mathcal{G_{HA}}$ is a sequence of locationslocations and edgesedges given as:
@@
where $l_{i}$ $\in$ LocLoc, $e_{i}$ $\in$ EdgeEdge, and $l_{i}$, $l_{i+1}$ are the source and destination of $e_{i}$ respectively. The length of a path is the number of edges it contains. $\Box$
@@
The graph of the hybrid automaton provides a higher abstraction of the domain in the sense that if there is no path from $l_{0}$ to $l_{goal}$ in $\mathcal{G_{HA}}$, then there cannot be any valid run of a plan from $Init$ to $Goal$ and hence the planning problem is unsolvable. We may then identify the cause of unsolvability to be in the discrete dynamics, oblivious to the continuous dynamics of the domain. More importantly, as we shall see now, a chain of sub-problems can be identified from the longest common subsequence of the strings in $PS(\Pi)$. Finding a longest common subsequence (LCS) between strings is a classic computer science problem. An LCS measures the closeness of two or more strings by finding the maximum number of identical symbols in them in the same order (DBLP:conf/spire/BergrothHR00; DBLP:journals/jacm/Maier78; princeton1974finding). Recall that a subsequence is different from a substring which additionally requires that the common symbols present in the strings are without gaps. We now present the main result of the paper.
@@
LCS of strings can be computed using the standard dynamic programming paradigm (DBLP:conf/spire/BergrothHR00). For instance, in the motivating example of planetary rover domain, the longest common sequence for the strings of paths of length bounded by 10 from $l_{11}$ to $l_{25}$ is $l_{11}$-$l_{6}$-$l_{1}$-$l_{2}$-$l_{3}$-$l_{8}$-$l_{13}$-$l_{14}$-$l_{25}$. In the next section, we show the computation of Explanation($\Pi$).